ganeshie8 (ganeshie8):

Find the last two digits of the perfect number \[\large 2^{19936}\left(2^{19937}-1\right)\]

OpenStudy (alexandervonhumboldt2):

hmm i did such stuff i need some time to think such stuff is in russians 8 grade which i'm studying now

OpenStudy (dan815):

bahaha

OpenStudy (alexandervonhumboldt2):

well 19936 and 19937 have difference of one i think this may be usful

OpenStudy (dan815):

u need to see the the last 2 digits of 2 to the 2 big numbers and 2 to the....36

ganeshie8 (ganeshie8):

Indeed thats a good observation! also 19937 is prime and this follows from the hypothesis that the number is a perfect number

OpenStudy (alexandervonhumboldt2):

it hought about trying to isolate 2^19366 as a sharable factor but it didn;t helped byffffff hmmm

ganeshie8 (ganeshie8):

idk if 19937 being a prime is useful in finidng the last two digits hmm

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (dan815):

2 goes in 1 2 4 8 16 32 64 28 56 12 24 48 96

OpenStudy (alexandervonhumboldt2):

2^19936 * 2^1 =2^19937 i'm trying to isolate

OpenStudy (freckles):

I always forget. Does last two mean the one's and ten's digit or the other way around?

OpenStudy (dan815):

10s and 1s

ganeshie8 (ganeshie8):

yes ones and tens...

OpenStudy (freckles):

That is always confusing to me because I see them as the first digits of a number but I don't know anything. :p

ganeshie8 (ganeshie8):

123456 you always write 56 last :P

OpenStudy (alexandervonhumboldt2):

maybe distribute

OpenStudy (freckles):

not when I add numbers or subtract them :p

ganeshie8 (ganeshie8):

this looks hard.. im trying to find the units digit for now :/

OpenStudy (alexandervonhumboldt2):

2^39873 - 2^19936 if 2 is in power of 2 4 , 6 , 8, 10 and ........... the numebr will be also умут

OpenStudy (alexandervonhumboldt2):

hmm

OpenStudy (freckles):

will n mod 10 always give you the 1's digit of n? and n mod 100 always give you 1's and 10's digit of n?

ganeshie8 (ganeshie8):

yes n mod 10 gives the units digit of n

OpenStudy (dan815):

i found theres a repetition in digits 2

OpenStudy (alexandervonhumboldt2):

there is a repetition

OpenStudy (alexandervonhumboldt2):

crre t

ganeshie8 (ganeshie8):

repitition in powers of 2 ?

OpenStudy (dan815):

yeah every 20 numbers

ganeshie8 (ganeshie8):

\[2^{20+x}\pmod{100} = 2^x \pmod{100}\] ?

OpenStudy (dan815):

12 24 48 96 92 84 68 36 72 44 88 76 52 04 08 16 32 64 28 56 12 <-repeat 24 .

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (alexandervonhumboldt2):

every number repetas thata sequaence

ganeshie8 (ganeshie8):

thats an interesting pattern ! but how is it useful

OpenStudy (dan815):

>> mod((19936+19937),20) ans = 13 >> mod((19936),20) ans = 16

OpenStudy (dan815):

lets hope 3rrd digit is not needed

OpenStudy (alexandervonhumboldt2):

wee get 2 gihts and we need only 2 so this is useful

OpenStudy (dan815):

12 <----- 2^1 (techically 9 but...) 24 48 96 92 84 68 36 72 44 88 76 52 04 08 16 32 64 28 56 12 <-repeat

OpenStudy (dan815):

52 - 16

OpenStudy (dan815):

36

OpenStudy (dan815):

are the last 2 digits 36?

ganeshie8 (ganeshie8):

Very close! answer is 56

OpenStudy (alexandervonhumboldt2):

56! it is in our secuane!!!

OpenStudy (dan815):

aye back to the drawing board

OpenStudy (dan815):

oh oops a mistake!

OpenStudy (dan815):

2^13= 8192 2^16=...36 soo its 56

OpenStudy (dan815):

i forgot that thing was start at 2^9 anot not 2^1

OpenStudy (dan815):

http://prntscr.com/6dyroc

OpenStudy (alexandervonhumboldt2):

@dan815 yeah you are correct i thought it is much easier but i havent learned that stuff yer. we have been solving easier.

OpenStudy (dan815):

haha alexander are u in russia?

OpenStudy (alexandervonhumboldt2):

well no how not in russia but i'm studying suuing russian materials and books

ganeshie8 (ganeshie8):

wow! looks pretty neat.. how did you work out 13 and 16 ?

OpenStudy (dan815):

> mod((19936+19937),20) ans = 13 >> mod((19936),20) ans = 16 >> 2^13

ganeshie8 (ganeshie8):

i mean how do you know the last two digits of given number will be same as the last two digits of 2^13 - 2^16

OpenStudy (dan815):

so we check what 2^13 is and 2^16

OpenStudy (dan815):

oh because

OpenStudy (dan815):

u got 2^x(2^y-1) so 2^(x+y)-2^x = .........92 -.......36 --------- you just hope the 2nd digit on the bottom wasnt higher than the top

OpenStudy (dan815):

and in our case it wasnt so no need for 3rd digit

OpenStudy (alexandervonhumboldt2):

92-36=56 so yeah

ganeshie8 (ganeshie8):

how did 19936+19937 became 13 ?

ganeshie8 (ganeshie8):

nvm i see XD

OpenStudy (dan815):

something to look into tho.... 36 was in the 20 number seequene and so was 56

OpenStudy (dan815):

is it possible that 2^n*(2^(n+1)-1) is always in the sequence of 20 digits

OpenStudy (dan815):

ah nvm thers a fail at 2^7(2^8-1)

ganeshie8 (ganeshie8):

the last two digits must be in the sequence of those 20 numbers whenever n+1 is prime

OpenStudy (dan815):

oh really lemme check some

OpenStudy (dan815):

wow thats cool

OpenStudy (dan815):

why is that happening?

ganeshie8 (ganeshie8):

actually it might not work always :/ one sec..

ganeshie8 (ganeshie8):

we know these : 1) if \(2^{n-1}(2^n-1)\) is a perfect number then \(n\) is prime 2) if \(2^n-1\) is a prime number then \(n\) is prime

ganeshie8 (ganeshie8):

lets see if those remainders really fall in the powers of 2 : when \(n\) is prime, \(n\) will be in one of two forms : 4k+1, 4k+3

OpenStudy (dan815):

what is the def of perfect number

ganeshie8 (ganeshie8):

6 is perfect number because 6 = 1+2+3 a number which equals the sum of its proper divisors is called a perfect number

ganeshie8 (ganeshie8):

Claim 1 : if \(n\) is in form \(4k+3\) then the last two digits of perfect number \(2^{n-1}(2^n-1)\) is always \(28\)

ganeshie8 (ganeshie8):

6 just happens to have first 3 natural numbers as its divosors

OpenStudy (dan815):

ah ya sry

ganeshie8 (ganeshie8):

but there is something related to what you found just now : every perfect number is a triangular number

OpenStudy (dan815):

if 2^n-1 is a prime then n is prime hmm lemme think about this for a sec

ganeshie8 (ganeshie8):

kk

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \normalsize \text{by euler totient function}\hspace{.33em}\\~\\ \phi{100}=40 \hspace{.33em}\\~\\ \dfrac{2^{19936}\left(2^{19937}-1\right)}{100}\hspace{.33em}\\~\\ \implies \dfrac{2^{498\times 40+16}\left(2^{498\times 40+17}-1\right)}{100}\hspace{.33em}\\~\\ \implies \dfrac{2^{16}\left(2^{17}-1\right)}{100}\hspace{.33em}\\~\\ \implies \dfrac{2^{16}\left(2^{17}-1\right)}{100}\hspace{.33em}\\~\\ \color{red}{\implies} \dfrac{2^{16}\left(2^{17}-1\right)}{100}\hspace{.33em}\\~\\ \end{align}}\) this minimizes

ganeshie8 (ganeshie8):

that looks neat !!!

OpenStudy (dan815):

wow..

OpenStudy (dan815):

i dont know how that happened but wow..

OpenStudy (dan815):

euler was a cool guy

OpenStudy (mathmath333):

i m thinking that their can be much short method with binomial somehow

OpenStudy (dan815):

ya i was thinking about binomials too

OpenStudy (dan815):

there might be a simple method to find the 2nd digit with binomials expansions

OpenStudy (mathmath333):

yes binomial is very effective particularly in cases in \(\pmod {100}\) and \(\pmod {10}\)

OpenStudy (dan815):

oky i think i see it because if n is not prime then it will include a composition of primes other than 2

OpenStudy (dan815):

which cannot exist sinice its 2^p=2*n+2 and if n is a composition of 2s that is also not possible since ull have 2^P=2^k *(some product of primes)+2

OpenStudy (dan815):

wich simply cannt be since u will have 2^P = 2*(2^(k-1)*(product of primes)/2 +1) ^ not possible of the product doesnt exist then its not possible because of the1

OpenStudy (dan815):

because that product of primes cannot include a 2

ganeshie8 (ganeshie8):

\(2^n - 1\) need not be a prime when \(n\) is prime

OpenStudy (dan815):

waiit whatt

ganeshie8 (ganeshie8):

\[2^n-1 \text{ is prime } \implies n \text{ is prime}\] the converse wont hold

OpenStudy (dan815):

arghh i see okay

OpenStudy (dan815):

show me the proof for that statement 2)

OpenStudy (dan815):

ill check in a bit i got ppl overright now

ganeshie8 (ganeshie8):

which one

OpenStudy (dan815):

why 2^n-1 if its prime then n must be prime